Aki's bicycle designs has determined that when x hundred bicycles are built the average cost per bicycle is given by C(x)=0.1x^2-1.1x+7.715 where C(x) is in hundreds of dollars. How many bicycles should the shop build to minimize the average cost per bbicycle?
Answer by nerdybill(7384)

Aki's bicycle designs has determined that when x hundred bicycles are built the average cost per bicycle is given by C(x)=0.1x^2-1.1x+7.715 where C(x) is in hundreds of dollars. How many bicycles should the shop build to minimize the average cost per bbicycle?
.
We know this is a parabola that opens upwards (by looking at the coefficient associated with the x^2 term) -- that is, the vertex will be the minimum.
.
x = -b/(2a)
x = -(-1.1)/(2*0.1)
x = (1.1)/(0.2)
x = 5.5
since x was in terms on hundreds:
number of bicycles to minimize ave cost is
5.5*100 = 550